OpenStudy (anonymous):

Pretest for school. No grade just need a better understanding as well as a solution? I want to be prepared and have a good piece of mind walking into school monday. "Simplify by algebraic ratios: x^4 - 3x^3 over x^2 - 9

ganeshie8 (ganeshie8):

hint- x^2-9 = x^2-3^2 ( you can factor this using the identity a^2-b^2 = (a+b)(a-b) )

ganeshie8 (ganeshie8):

does that help.... you can finish the problem wid that hint ?

OpenStudy (anonymous):

I still can't do the bottom one. I have never learned this before. It's a semester pretest.

ganeshie8 (ganeshie8):

ohhk... its easy. bottom can be written like this : x^2-9 = x^2-3^2 = (x+3)(x-3)

OpenStudy (anonymous):

Oh :D thanks that helped a lot

ganeshie8 (ganeshie8):

glad to hear that ! (: top you can take x^3 common, right ?

OpenStudy (anonymous):

yes so it's (x-3)(x+3) over (x+3)(x-3)

ganeshie8 (ganeshie8):

\(\huge \frac{x^4-3x^3}{(x+3)(x-3)}\)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

I was still off haha

ganeshie8 (ganeshie8):

thats the given expression.... we just factored the bottom part.

ganeshie8 (ganeshie8):

lets factor top... okay ?

ganeshie8 (ganeshie8):

and see if something cancels

ganeshie8 (ganeshie8):

\(\huge \frac{x^4-3x^3}{(x+3)(x-3)}\) \(\huge \frac{x^3(x-3)}{(x+3)(x-3)}\) \(\huge \frac{x^3\cancel{(x-3)}}{(x+3)\cancel{(x-3)}}\) \(\huge \frac{x^3}{(x+3)}\)

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

complete solution there. see if it makes sense....

OpenStudy (anonymous):

Way better understanding of it now.

ganeshie8 (ganeshie8):

great ! yw.. .. . :D